OpenStudy (anonymous):

what are the steps of a finding the area of a garden that is 3 yards by 4 feet please help!!!

OpenStudy (anonymous):

Area is found by the formula A=b*h b= base h=height A=area

OpenStudy (anonymous):

remember that you must have similar units (1yd = 3 ft)

OpenStudy (anonymous):

would that be all ?

OpenStudy (anonymous):

Yes, it's a relatively straightforward problem. Convert the units and use the formula

OpenStudy (anonymous):

yes.... finding area is finding the number of squares that fit in a region....

OpenStudy (anonymous):

i dont really know how to convert the units

OpenStudy (anonymous):

I'll show you an example. Coverting inches to feet. 12in = ft, convert 36 inches to feet \[36in\frac{1ft}{12in}\frac{36in*1ft}{12in}=3*1ft=3ft\] Notice how the inchest cancel each other out in the conversion.

OpenStudy (anonymous):

Do the same from yard to feet, 1yd = 3ft

OpenStudy (anonymous):

so i would convert 3 yards to feet

OpenStudy (anonymous):

yes to get the same units as the 4ft

OpenStudy (anonymous):

i kind of understand it

OpenStudy (anonymous):

work it out and i'll walk you through it

OpenStudy (anonymous):